# MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression

### v2

@@ Abstract
footnotetext: Corresponding author: Xuefei Ning (foxdoraame@gmail.com)
@@ 1 Introduction
Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval benchmark (Li et al., 2023). This retrieval benchmark takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions. (a) Original model with a full attention span; (b) StreamingLLM with half the attention span, showing reduced effectiveness beyond the span; (c) MoA with half the attention span, maintaining effectiveness beyond the span.
@@
Large Language Models (LLMs) exhibit remarkable versatility across numerous applications (Brown et al., 2020; Tay et al., 2022; Wan et al., 2023). Central to LLM is the attention mechanism (Vaswani et al., 2017), which computes interactions among tokens within a certain span, thereby enabling context understanding. Scaling input length is crucial for enhancing LLM capabilities (Chen et al., 2023; Tworkowski et al., 2023), including fact retrieval, summarization, few-shot learning, question answering and so on (Bai et al., 2023; Yuan et al., 2024). However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges (Sheng et al., 2023; Xiao et al., 2024c; Han et al., 2023; Kwon et al., 2023).
@@
Previous work proposes sparse attention methods to address the efficiency challenges of long contexts in generative LLMs. These methods typically employ a uniform, fixed-span sliding window mask across all heads and input lengths, limiting attention to local contexts only (Xiao et al., 2024c; Han et al., 2023). This approach allows the LLM to take long inputs with a fixed attention span, keeping bounded attention computation and KV caching overhead. Following previous works (Chen et al., 2023; Tworkowski et al., 2023), we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold. In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span (Feng et al., 2022; Zaheer et al., 2020). Nonetheless, we reveal that uniform masks, like StreamingLLM (Xiao et al., 2024c), hardly extend effective context length beyond the span, as shown in Figure 6. Figure 1(b) further illustrates such limitation: with a 50% attention span mask, StreamingLLM fails to accurately retrieve content from the earlier half of the input and performs even worse at longer input lengths. Figure 2 reveals one possible explanation for the problem: while some attention heads focus on local contexts, others encompass the broad span of the entire input sequence. Consequently, the uniform approach fails to achieve a long effective context length as it limits the attention span of the global-context heads, while excessively allocates compute and memory budget for local-context heads. Additionally, as the input length increases, some attention heads need a faster increase in attention span than others to avoid serious performance degradation, as shown in Table 1. Unfortunately, the uniform approaches do not include heterogeneous rules to scale the attention spans differently for various heads. Besides, existing model compression methods (Men et al., 2024; Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024a; Kim et al., 2023; Li et al., 2024b) use general language modeling corpora to decide the compression plan, which cannot accurately profile the influence of compression on long-context tasks.
@@
In this work, we propose Mixture of Attention (MoA), a training-free sparse attention method. As illustrated in Figure 3, MoA constructs the search space of heterogeneous elastic rules of attention spans. For automatic LLM compression, MoA first utilizes gradient-based profiling to inspect the influences of each attention position on the prediction loss. Based on the profiling results, MoA tailors heterogeneous sparse attention configurations for each model layer and attention head. During profiling, MoA employs a calibration dataset with long-range dependencies and uses the original dense model’s response instead of the human-written response as the reference to calculate the loss. This ensures an accurate profiling of the attention influences to facilitate better compression results. Our contributions are summarized as follows.
@@
  • Heterogeneous Elastic Rules. We propose heterogeneous elastic rules for masks of each attention head. We formulate MoA compression search space to include a diverse range of elastic rules that tailor the local attention span relative to the input length for each attention head. The heterogeneous elastic rules improve the fact retrieval accuracy of MoA from 25% to 98% compared with masks with uniform span and scaling function for each head.
@@
  • Automatic Optimization. We propose an automatic pipeline to find the optimal compression plan encompassing heterogeneous elastic rules for various attention heads. This pipeline can efficiently find the optimal plan within several hours, for example, two hours for compressing Vicuna-13B.
Experiments show that MoA achieves $\times6$ 8.2 × throughput improvements over FlashAttention2, 1.71.9×1.7-1.9$\times$1.7 - 1.9 × over vLLM framework on 7B and 13B dense LLMs at a 50% density (the average of KV-Cache length / input length), with only a 1% average relative degradation in retrieval accuracy. In Section 6.4, we show that this significant throughput improvements of MoA over FlashAttention2 can be attributed to four factors: (1) the static size of the KV-cache, (2) reduced attention computations, (3) increased batch size enabled by reduced memory usage, and (4) a specialized kernel implementation. Additionally, MoA achieves over 90% retrieval accuracy with just 25% average density, far surpassing sparse attention baselines that need a density of 75% to 100% for similar performance. On long-context understanding benchmarks, MoA performs comparably to dense models, with a maximum relative performance drop of less than 5%, which is about one-sixth of that observed with the uniform sparse attention baseline. Our code is available at https://github.com/thu-nics/MoA.
@@ 2.1 Attention mechanism
The Multi-Head Self Attention (MHA) mechanism (Vaswani et al., 2017) is crucial to the functionality of LLMs. It starts with an input sequence transformed into query (Q), key (K), and value (V) matrices through linear projections. These matrices, combined with the cached K and V (KV-Cache) from previous sequences, compute the attention matrix (A). This calculation is modified by a causal mask (M) to ensure autoregressive properties, resulting in the output (O), as depicted in Equation 1:
@@
Autoregressive inference in LLMs involves two stages: prefill and decode. During prefill, the model processes the entire input sequence to generate the initial response token. In the subsequent decode stage, it uses the newly generated token and previously cached K and V matrices to produce subsequent tokens until the generation concludes. Although effective, this iterative process increases memory and computation demands due to the expanding KV-Cache.
2.2 Efficient Attention
@@
Efficient methods are proposed to mitigate the computation and memory costs associated with attention. One branch of work uses dynamic sparse attention masks to adaptively skip attention computations during prefill stage (Pagliardini et al., 2023; Qu et al., 2022; Roy et al., 2021; Wang et al., 2021; Lu et al., 2021; Kitaev et al., 2020) or drop KV-Cache during decode stage (Anagnostidis et al., 2023; Zhang et al., 2023; Ge et al., 2023; Sheng et al., 2023; Liu et al., 2023a) based on the input sequences. However, due to the complex control and computation flow, dynamic prefill often requires specific hardware to achieve substantial wall-time speedup (Qu et al., 2022; Wang et al., 2021; Lu et al., 2021; Ham et al., 2021; 2020). Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining (Anagnostidis et al., 2023), additional accumulated attention score computation (Sheng et al., 2023; Zhang et al., 2023; Liu et al., 2023a; Ge et al., 2023), or extensive memory swap for KV-Cache retrieval (Tang et al., 2024; Xiao et al., 2024a).
@@
Another branch of work uses static sparse attention, where predefined masks are applied consistently across all processed sentences. Thanks to the fixed computation flow, static sparse attention is generally more efficient and GPU-friendly. For language understanding models such as BERT (Devlin et al., 2018), various masks are used (Zaheer et al., 2020; Beltagy et al., 2020; Child et al., 2019; Zhou et al., 2024; Xiao et al., 2024c; Han et al., 2023). But for generative LLMs, the predominant method is the fixed-span sliding window mask with global attention on a few initial tokens (Xiao et al., 2024c; Han et al., 2023). With the local attention pattern, the KV-Cache beyond the current attention span can be dropped, saving much memory for long sequence scenarios. However, the uniform static masks across different attention heads and input lengths are model- and data-agnostic, which can compromise LLMs’ effective context length and lead to suboptimal performance in long sequence scenarios. Our method falls within this category, benefiting from the efficiency and training-free advantages, while addressing the performance limitations encountered by previous methods.
In addition to sparse attention, alternative mechanisms have been proposed to replace traditional attention for long-sequence modeling (Gu & Dao, 2023; Peng et al., 2023; Sun et al., 2023; Poli et al., 2023; Li et al., 2022; Kacham et al., 2023; Peng et al., 2021; Choromanski et al., 2020; Wang et al., 2020). However, these new mechanisms often require different weights compared to vanilla transformers, imposing significant re-training overhead for LLMs.
@@
Previous works also propose LLM acceleration frameworks (Gugger et al., 2022; Aminabadi et al., 2022; Sheng et al., 2023; Kwon et al., 2023), as well as kernel-level optimizations (Dao et al., 2022; Dao, 2023; Shah et al., 2024). These kernel and system optimizations are orthogonal to our work and can be integrated to further enhance efficiency.
3 Mixture of Attention (MoA)
@@
We first illustrate the heterogeneity of the attention patterns in pre-trained LLMs in Section 3.1. Based on this insight, we define the search space for our Mixture-of-Attention (MoA) method in Section 3.2.
@@
Table 1: Retrieval accuracy of Vicuna-7B with sliding-window sparse attention across various model layers, window spans, and input lengths.
@@ 3.1 Mixture of Attention Patterns and Elastic Rules
Heterogeneous Attention Patterns. Different attention heads in LLMs exhibit heterogeneous attention patterns, as shown in Figure 2. For example, the first head primarily focuses on local contexts with a narrow-span sliding window, while the third head covers nearly the entire input, indicating global attention. The attention spans of different heads mostly remain constant across various tasks and datasets, as shown in Appendix D.1. Table 1 demonstrates that applying the same sliding-window sparse attention mask across model layers can lead to a 65% variance in retrieval accuracies. It conforms to the multi-head self-attention design principle of capturing varied information (Vaswani et al., 2017), as well as the findings from concurrent research that identifies specific attention heads for global text retrieval (Wu et al., 2024).
Heterogeneous Elastic Rules. In addition to heterogeneity at a certain length, different attention heads also exhibit varying elastic behaviors as the input length changes. Figure 2 illustrates this variability: for shorter inputs (the upper left part of the attention matrix), the second and third heads initially show global attention. However, as input length increases, the second head remains the medium-span local focus, while the third head continues to expand as global attention. Table 1 further evidences the diverse elastic rules. For example, at 4k input length, a 2k sliding-window sparse attention mask on layers 9 to 11 yields better retrieval accuracy than on layers 17 to 19. However, the opposite is true for an 8k input length. This data supports the visual observations from Figure 2, highlighting that attention patterns respond to input length scaling differently. Leveraging these insights, MoA encompasses heterogeneous elastic rules as the search space.
@@ 3.2 Heterogeneous Elastic Rule Search Space
In designing the search space for the MoA mask, we consider the inherently heterogeneous and elastic nature of LLM attention patterns. As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base sparse attention mask (Beltagy et al., 2020). Following previous work (Xiao et al., 2024c; Han et al., 2023), the initial few tokens (64 tokens for MoA) are not masked. The attention span equals the sliding-window-span plus the number of initial unmasked tokens. We define the attention span S of head $h$ at input length N using a straightforward linear function:
@@
where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length of a specific attention head.
The $\alpha$ and $\beta$ hyperparameters of each attention head can be chosen from multiple discrete options. For LLMs with many layers, the MoA search space can become quite large. For example, for a 7B model consisting of $32$ attention heads and $32$ layers, with $6$ and $9$ different $\alpha$ and $\beta$ options for each head, the potential search space expands to $54^{1024}$ configurations. Thus, we design the automatic pipeline to efficiently pinpoint the optimal $\alpha$s and $\beta$s for any LLM.
4 Automatic Pipeline for MoA Compression
Figure 3: Overview of the MoA. (a) The sparse attention search space includes heterogeneous elastic rules of the attention span on sliding-window masks. (b) The automatic compression pipeline begins with a calibration dataset, which includes long-dependency contexts and supervision texts generated by the original dense LLM. MoA profiles each attention value’s impact on model predictions within this dataset, revealing accuracy losses for different candidate elastic rules across various input lengths. The final optimization step selects elastic rules for each attention head to minimize the total prediction loss while adhering to specified density constraints.
This section outlines the MoA automatic compression pipeline as shown in Figure 3(b). Starting with a trained LLM and a calibration dataset, MoA first profiles the influence of each attention value on the model’s prediction loss for various input sequences from the calibration dataset. The masked sum of the influences represents the accuracy loss associated with each mask at different input lengths, showing the accuracy loss each candidate elastic rule could cause at that length. Then, MoA optimizes the compression plan by selecting the optimal elastic rule for each head, which minimizes the accuracy loss across various lengths while adhering to specified density constraints. The following sections provide detailed discussions of each step in this pipeline.
@@ 4.1 Attention Influence Profiling
In the profile step, MoA quantifies the impact of individual attention values on the final prediction loss of a pre-trained LLM. It informs the subsequent step about the influence of masking each attention value, revealing the accuracy trade-offs of the candidate elastic rules for each attention head.
The influence of each attention value is derived from the attention matrix A and its gradient $\partial L/\partial\textbf{A}$, computed over a calibration dataset. When applying sparse attention masks, we approximate the change in the model’s prediction loss, $\Delta L$, using a first-order Taylor expansion based on variations in the attention matrices A: $\Delta L=\sum_{h}\sum_{i}\sum_{j}\partial L/\partial A_{h,i,j}\cdot\Delta A_{h,i,j}$. Here, $h$ indexes the attention heads across all layers, and $i,j$ are the row and column indices within each attention matrix $\textbf{A}_{h}$.
We define the attention influence matrix, $\textit{E}_{h,i,j}$, as the estimated change in loss, $\Delta L$, if the attention value $A_{h,i,j}$ is masked (i.e., set to zero). As shown in Equation 3, this measure considers both the direct and indirect effects of the mask. For notation simplicity, we omit the head index $h$ here. Initially, masking directly reduces the attention value to zero, represented by $\Delta A_{i,j|j}=-A_{i,j}$. Additionally, the softmax function in attention normalizes the sum of each row in the attention matrix to one. Thus, setting one attention value at column $j$ to zero causes an increase in the other attention values, $\Delta A_{i,n|j},n\neq j$, within the same row. These two effects are integrated into the following formulation, whose derivation is provided in Appendix D.2:
@@
In practice, we use backpropagation on a calibration dataset to calculate the average attention influence $\bar{\textbf{E}}_{h}$ of each head across data items. The average attention influence is calculated respectively for different input lengths. The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative in deep learning frameworks like PyTorch (Paszke et al., 2019). The detailed calibration dataset setup is discussed in Section 5.
With the average attention influence of each head, MoA can calculate the accuracy loss of applying a candidate elastic rule at a specific input length. The loss is calculated with the sum of masked average attention influence according to the rule. We denote $\textbf{M}_{r_{h}}$ as the binary mask at head $h$ that corresponds to rule $r$, with masked positions marked as 1 and others as 0. We formalize accuracy loss $\Delta L$ as follows:
@@
After the profile stage, MoA acquires the unique accuracy-density trade-offs of elastic rules. It informs the allocation of denser masks to more sensitive heads and lighter masks to less sensitive ones. Profiling at different input lengths enables the identification of the most effective elastic rules, even for unseen lengths.

### v3

# Mixture of Attention Spans: Optimizing LLM Inference Efficiency with Heterogeneous Sliding-Window Lengths

@@ 1 Introduction
Large Language Models (LLMs) exhibit remarkable versatility across numerous applications (brown2020gpt3; tay2022efficientSurvey; Wan2023EfficientLLM). Central to LLM is the attention mechanism (vaswani2017attention), which computes interactions among tokens within a certain span, thereby enabling context understanding. Scaling input length is crucial for enhancing LLM capabilities (Chen2023PI; Tworkowski2023FocusedTrans), including fact retrieval, summarization, few-shot learning, question answering and so on (bai2023longbench; yuan2024lveval). However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges (xiao2023streamingLLM).
Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval benchmark (lmsys2023longeval). The benchmark takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions. (a) Original model with a full attention span; (b) StreamingLLM with half the attention span, showing reduced effectiveness beyond the span; (c) MoA with half the attention span, maintaining effectiveness beyond the span.
@@
Previous work has proposed sliding-window attention to address the efficiency challenges of long contexts in generative LLMs. These methods typically employ a uniform, fixed-span sliding-window mask across all heads and input lengths, limiting attention to local span only (xiao2023streamingLLM; han2023lmInfinite). This design allows the LLM to process long inputs within a bounded attention computation and KV caching overhead. Following previous works (Chen2023PI; Tworkowski2023FocusedTrans), we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold. In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span (feng2022diffuser; zaheer2020bigbird). Nonetheless, we reveal that uniform sliding-window methods, such as StreamingLLM (xiao2023streamingLLM), hardly extend effective context length beyond the span. As shown in Figure 1(b), with a 50% attention span, StreamingLLM fails to accurately retrieve earlier half of the input and performs even worse at longer input lengths. Figure 2 reveals one possible explanation for the problem: while some attention heads focus on local contexts, others encompass the broad span of the entire input sequence. Consequently, the uniform approach fails to achieve a long effective context length, as it limits the attention span of global-context heads while excessively allocating compute and memory budgets to local-context heads. Additionally, as the input length increases, some attention heads need a faster increase in attention span than others to avoid serious performance degradation, as shown in Table 1. Unfortunately, the uniform approaches do not include heterogeneous rules to scale the attention spans differently for various heads. In addition, existing model compression methods, including quantization and sparse attention (smooth_quant; llm-mq; squeezellm; li2024evaluating), typically rely on general language modeling corpora to determine compression configurations, which fail to accurately capture their impact on long-context tasks.
@@
In this work, we propose Mixture of Attention Spans (MoA), a training-free heterogeneous sliding-window attention method. As shown in Figure 3, MoA constructs a search space of heterogeneous and elastic rules that scale each head’s window length with input length. To automate rule selection, MoA first utilizes gradient-based profiling to inspect the influences of each attention position on the prediction loss. Based on the profiling results, MoA tailors heterogeneous window length for each attention head and input length. During profiling, MoA employs a calibration dataset with long-range summaries from a full-attention model, instead of the human-written summaries, as the reference to calculate the loss. This ensures an accurate profiling of attention influences to facilitate better performance. Our contributions are summarized as follows.
@@
  • Heterogeneous Elastic Rules. We propose heterogeneous elastic rules for sliding-window length of each attention head. We formulate MoA configuration search space to include a diverse range of elastic rules that tailor the local attention span relative to the input length for each attention head. The heterogeneous elastic rules improve content retrieval accuracy from 25% to 98% over the uniform counterpart.
@@
  • Automatic Configuration Search. We propose an automatic pipeline to find the optimal configuration encompassing heterogeneous elastic rules for various attention heads. This pipeline efficiently finds the optimal configuration within several hours—for example, two hours for Vicuna-13B.
[…]
@@ 2.1 Attention Mechanism
The Multi-Head Self-Attention (MHA) mechanism (vaswani2017attention) is crucial to the functionality of LLMs. It starts with an input sequence transformed into query (Q), key (K), and value (V) matrices through linear projections. These matrices, combined with the cached K and V (KV-Cache) from previous sequences, compute the attention matrix (A). This calculation is modified by a causal mask (M) to ensure autoregressive properties, resulting in the output (O), as depicted in Equation 1:
@@
Autoregressive inference in LLMs involves two stages: prefill and decode. During prefill, the model processes the entire input sequence to generate the initial response token. In the subsequent decode stage, it iteratively uses the newly generated token and previously cached K, V matrices to produce subsequent tokens. Although effective, this iterative process increases memory and computation demands due to the expanding KV-Cache.
2.2 Efficient Attention Methods
@@
Efficient attention methods have been proposed to mitigate the computation and memory costs. One branch of work dynamically skip attention computations during the prefill stage (Pagliardini2023SparseFlash; qu2022dota; roy2021routingTransformer; wang2021spatten; lu2021sanger; kitaev2020reformer) or drop the KV-Cache during the decode stage (Anagnostidis2023DynamicPruning; Zhang2023H2O; Ge2023FastGen; sheng2023flexgen; Liu2023Scissorhands) based on the input sequence. However, due to the complex control and computation flow, dynamic sparse prefill often requires specific hardware to achieve substantial wall-time speedups (qu2022dota; wang2021spatten; lu2021sanger; ham2021elsa; ham2020A3). Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining (Anagnostidis2023DynamicPruning), additional pruning score computation (sheng2023flexgen; Zhang2023H2O; Liu2023Scissorhands; Ge2023FastGen; li2024snapkv; cai2024pyramidkv), or extensive memory swapping for KV-Cache retrieval (tang2024quest; xiao2024infllm).
@@
Another branch of work uses static sparse attention, where predefined masks are applied consistently across all input sequences. Thanks to the fixed computation flow, static sparse attention is generally more efficient and GPU-friendly. For language understanding models such as BERT (devlin2018bert), various masks are used (zaheer2020bigbird; beltagy2020longformer; child2019spTrans; Zhou2024EfficientSurvey; xiao2023streamingLLM; han2023lmInfinite). But for generative LLMs, the predominant method is the uniform sliding-window mask with global attention on a few initial tokens (xiao2023streamingLLM; han2023lmInfinite). For these approaches, the KV-Cache beyond the local attention span can be dropped, saving much memory for long sequence scenarios. However, uniform sliding-window masks across different attention heads and input lengths are model- and data-agnostic, which can compromise LLMs’ effective context length and lead to suboptimal performance in long-sequence scenarios. Our method falls within this category, benefiting from the efficiency and training-free advantages, while addressing the performance limitations encountered by previous methods.
@@
Previous works also propose LLM acceleration frameworks (huggingface2022accelerate; Aminabadi2022DeepSpeed; sheng2023flexgen; Kwon2023vllm), as well as kernel-level optimizations (dao2022flashattention; dao2023flashattention2; shah2024flashattention3). These kernel and system optimizations are orthogonal to our work and can be integrated to further enhance efficiency.
3 Mixture of Attention Spans (MoA)
@@
We first illustrate the heterogeneity of attention patterns in pre-trained LLMs in Section 3.1. Based on this insight, we define the search space for our MoA method in Section 3.2.
@@
Table 1: Retrieval accuracy of Vicuna-7B with sliding-window attention across various model layers, window spans, and input lengths.
@@ 3.1 Mixture of Attention Patterns and Elastic Rules
Heterogeneous Attention Patterns. Different attention heads in LLMs exhibit heterogeneous attention patterns, as shown in Figure 2. For example, the first head primarily focuses on local contexts with a narrow-span sliding window, while the third head covers nearly the entire input, indicating global attention. The attention spans of different heads mostly remain constant across various tasks and datasets, as shown in Appendix E.1. Table 1 demonstrates that applying the same sliding-window attention mask across model layers can lead to a 65% variance in retrieval accuracies. This aligns with the multi-head self-attention design principle of capturing varied information (vaswani2017attention), as well as findings from concurrent research that identify specific attention heads for global text retrieval (wu2024retrievalHead).
Heterogeneous Elastic Rules. In addition to heterogeneity at a certain length, different attention heads also exhibit varying elastic behaviors as the input length changes. Figure 2 illustrates this variability: for shorter inputs (the upper-left part of the attention matrix), the second and third heads initially show global attention. However, as input length increases, the second head maintains a medium-span local focus, while the third head continues to expand as global attention. Table 1 further evidences the diverse elastic rules. For example, at a 4k input length, a 2k sliding-window attention mask on layers 9 to 11 yields better retrieval accuracy than on layers 17 to 19; however, the opposite holds for an 8k input length. These data support the visual observations from Figure 2, highlighting that attention patterns respond differently to input-length scaling. Leveraging these insights, MoA encompasses heterogeneous elastic rules as the search space.
@@ 3.2 Heterogeneous Elastic Rule Search Space
In designing the search space for MoA, we consider the inherently heterogeneous and elastic nature of LLM attention patterns. As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base attention mask (beltagy2020longformer). Following previous work (xiao2023streamingLLM; han2023lmInfinite), the initial few tokens (64 tokens for MoA) are not masked. The attention span equals the sliding-window span plus the number of initially unmasked tokens. We define the attention span S of head $h$ at input length N using a straightforward linear function:
@@
where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length for a specific attention head.
[…]
4 Automatic MoA Configuration Search Pipeline
Figure 3: Overview of MoA. (a) The search space includes heterogeneous elastic rules of the attention span on sliding-window masks. (b) The automatic rule search pipeline begins with a calibration dataset, which includes long-dependency contexts and supervision texts generated by the original dense LLM. MoA profiles each attention value’s impact on model predictions within this dataset, revealing accuracy losses for different candidate elastic rules across various input lengths. The final optimization step selects elastic rules for each attention head to minimize the total prediction loss while adhering to specified density constraints.
This section outlines the MoA automatic configuration search pipeline, as shown in Figure 3(b). Starting with a trained LLM and a calibration dataset, MoA first profiles the influence of each attention value on the model’s prediction loss for various input sequences from the calibration dataset. The masked sum of the influences represents the accuracy loss associated with each mask at different input lengths, showing the accuracy loss each candidate elastic rule could cause at that length. Then, MoA optimizes the window-lengths by selecting the optimal elastic rule for each head, which minimizes the accuracy loss across various input lengths while adhering to specified density constraints. The following sections provide detailed discussions of each step in this pipeline. Detailed definitions of key notions are in Appendix A. The user-friendly code interface is described in Appendix E.4.
@@ 4.1 Attention Influence Profiling
In the profiling step, MoA quantifies the impact of individual attention values on the final prediction loss of a pre-trained LLM. It informs the subsequent step about the influence of masking each attention value, revealing the accuracy trade-offs of the candidate elastic rules for each attention head.
The influence of each attention value is derived from the attention matrix A and its gradient $\partial L/\partial\textbf{A}$, computed over a calibration dataset. When applying sliding-window attention masks, we approximate the change in the model’s prediction loss, $\Delta L$, using a first-order Taylor expansion based on variations in the attention matrices A: $\Delta L=\sum_{h}\sum_{i}\sum_{j}\partial L/\partial A_{h,i,j}\cdot\Delta A_{h,i,j}$. Here, $h$ indexes the attention heads across all layers, and $i,j$ are the row and column indices within each attention matrix $\textbf{A}_{h}$. Details on the calibration dataset and the prediction loss $L$ are provided in Section 5.
We define the attention influence matrix, $\textit{E}_{h,i,j}$, as the estimated change in loss, $\Delta L$, if the attention value $A_{h,i,j}$ is masked (i.e., set to zero). As shown in Equation 3, this measure considers both the direct and indirect effects of the mask. For notation simplicity, we omit the head index $h$ here. Initially, masking directly reduces the attention value to zero, represented by $\Delta A_{i,j|j}=-A_{i,j}$. Additionally, the softmax function in attention normalizes the sum of each row in the attention matrix to one. Thus, setting one attention value at column $j$ to zero causes an increase in the other attention values, $\Delta A_{i,n|j},n\neq j$, within the same row. These two effects are integrated into the following formulation, whose derivation is provided in Appendix E.2:
@@
In practice, we use backpropagation on a calibration dataset to calculate the average attention influence $\bar{\textbf{E}}_{h}$ of each head across data items. The average attention influence is calculated respectively for different input lengths. The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative rule in deep learning frameworks like PyTorch (paszke2019pytorch). The detailed calibration dataset setup is discussed in Section 5.
With the average attention influence of each head, MoA can calculate the accuracy loss of applying a candidate elastic rule at a specific input length. The loss is calculated as the sum of masked average attention influence according to the rule. We denote $\textbf{M}_{r_{h}}$ as the binary mask at head $h$ that corresponds to rule $r$, with masked positions marked as 1 and others as 0. We formalize accuracy loss $\Delta L$ as follows:
@@
After the profiling stage, MoA acquires the estimated accuracy impact of elastic rules across attention heads. This informs the allocation of longer windows to more sensitive heads and shorter windows to less sensitive ones. Profiling at different input lengths enables the identification of the most effective elastic rules, even for unseen lengths.